# On local Galois deformation rings: generalised reductive groups

### v1

@@ 5.1. Definition of $X^{\mathrm{gen}}_{\operatorname{GL}_{d}}$ and $X^{\mathrm{gen},\tau}_{G}$
If $L^{\prime}$ is a finite extension of $L$ with then ring of integers $\mathcal{O}^{\prime}$ then it follows from the definition that $(X^{\mathrm{gen},\tau}_{G})_{\mathcal{O}^{\prime}}\cong X^{\mathrm{gen},\tau}_{G_{\mathcal{O}^{\prime}}}$. The main technical result of the paper is the computation of the dimension of $X^{\mathrm{gen},\tau}_{G}$. Since $\mathcal{O}^{\prime}$ is finite and flat over $\mathcal{O}$ the dimensions of $(X^{\mathrm{gen},\tau}_{G})_{\mathcal{O}^{\prime}}$ and $X^{\mathrm{gen},\tau}_{G}$ are the same and using Proposition 2.8 we may assume that $G$ is split, $G/G^{0}$ is constant and $G(\mathcal{O})\rightarrow(G/G^{0})(\mathcal{O})$ is surjective.
@@
Part (1) follows from Lemmas 4.15 and 5.5.
@@
If $\mathfrak{p}$ is the maximal ideal of $R^{\mathrm{ps}}_{\operatorname{GL}_{d}}$ then $\kappa(\mathfrak{p})=k$ and we equip the algebraic closure $\overline{\kappa(\mathfrak{p})}$ with the discrete topology. If $\mathfrak{p}\in P_{1}R^{\mathrm{ps}}_{\operatorname{GL}_{d}}$ then $\kappa(\mathfrak{p})$ is a local field and the valuation topology on $\kappa(\mathfrak{p})$ makes it into a topological $R^{\mathrm{ps}}_{\operatorname{GL}_{d}}$-algebra by [8, Lemma 3.17]. We equip the algebraic closure $\overline{\kappa(\mathfrak{p})}$ with the topology induced by a valuation, which extends the valuation on $\kappa(\mathfrak{p})$.
@@
Let $A^{\mathrm{gen}}_{\operatorname{GL}_{d}}\twoheadrightarrow A^{\mathrm{gen},\tau}_{G}$ be the quotient representing $X^{\mathrm{gen},\tau}_{G}$. We will denote by $\overline{X}^{\mathrm{gen},\tau}_{G}$ the special fibre of $X^{\mathrm{gen},\tau}_{G}$.
@@
Let $y:R^{\mathrm{ps}}_{\operatorname{GL}_{d}}\rightarrow\Omega$ be a ring homomorphism into an algebraically closed field $\Omega$ and and let $x\in X^{\mathrm{gen},\tau}_{G}(\Omega)$ be a point above $y$. Let $H$ be the smallest Zariski closed subgroup of $G_{\Omega}$ containing $\rho_{x}(\gamma_{1}),\ldots,\rho_{x}(\gamma_{N})$, where $\rho_{x}$ is the specialisation at $x$ of the universal representation $\rho:\Gamma\rightarrow\operatorname{GL}_{d}(A^{\mathrm{gen}}_{\operatorname{GL}_{d}})$. Then $H$ is equal to the Zariski closure of $\rho_{x}(\Gamma)$ in $G_{\Omega}$.
@@ 5.2. Completion at maximal ideals and deformation problems
Let $Y\subset X^{\mathrm{gen},\tau}_{G}$ be the preimage of the closed point of $X^{\mathrm{ps}}_{\operatorname{GL}_{d}}$ and let $x$ be either a closed point of $Y$ or a closed point of $X^{\mathrm{gen},\tau}_{G}\setminus Y$ and let $y$ be its image in $\operatorname{Spec}R^{\mathrm{ps}}_{\operatorname{GL}_{d}}$. It follows from [8, Lemmas 3.17, 3.18] that $\kappa(x)$ is a finite extension of $\kappa(y)$ and there are the following possibilities:
@@
If $\rho_{A}\in D^{\square}_{\rho,G}(A)$ then [8, Proposition 3.33] implies that $A$ is naturally an $R^{\mathrm{ps}}_{\operatorname{GL}_{d}}$-algebra and $\tau\circ\rho\in X^{\mathrm{gen}}_{\operatorname{GL}_{d}}(A)$. Lemma 5.5 implies that $\rho_{A}$ factors through the topologically finitely generated quotient $Q$ of $\Gamma$. Hence, we may assume without loss of generality that $\gamma_{1},\ldots,\gamma_{N}$ are topological generators of $\Gamma$.
@@
By passing to the limit and using Lemma 5.15 (in a noetherian setting) we obtain
@@
If $G=\operatorname{GL}_{d}$ and $\tau=\operatorname{id}$ then the proposition is [8, Proposition 3.33], so that $R^{\square}_{\tau\circ\rho_{x},\operatorname{GL}_{d}}$ is naturally isomorphic to the completion of the local ring of $\Lambda\otimes_{\mathcal{O}}A^{\mathrm{gen}}_{\operatorname{GL}_{d}}$ at $\mathfrak{q}$. Since $A^{\mathrm{gen},\tau}_{G}=A^{\mathrm{gen}}_{\operatorname{GL}_{d}}\otimes_{\mathcal{O}(\operatorname{GL}_{d}^{N})}\mathcal{O}(G^{N})=A^{\mathrm{gen}}_{\operatorname{GL}_{d}}/IA^{\mathrm{gen}}_{\operatorname{GL}_{d}}$ by definition, and $R^{\square}_{\tau\circ\rho_{x},\operatorname{GL}_{d}}$ is noetherian by Proposition 3.8, the assertion follows from Lemma 5.15 applied to $A=(\Lambda\otimes_{\mathcal{O}}A^{\mathrm{gen}}_{\operatorname{GL}_{d}})_{\mathfrak{q}}$. ∎
@@
The assertion follows from Proposition 5.16 and [8, Lemmas 3.36, 3.37]. ∎
@@ 6. GIT quotients
The closed immersion $\overline{X}^{\mathrm{gen},\tau}_{G}\hookrightarrow X^{\mathrm{gen},\tau}_{G}$ induces a morphism $\overline{X}^{\mathrm{git},\tau}_{G}\rightarrow(X^{\mathrm{git},\tau}_{G})_{k}$, which is finite by [52, Theorem 2] and induces a bijection on geometric points by [52, Theorem 3]. We deduce that this map is a homeomorphism.
@@
[30, Exp. V, Thm. 4.1 (ii)] applied with $X_{0}=H^{N}\sslash H^{0}$ and $X_{1}=(H/H^{0})\times_{S}X_{0}$, implies that the morphism $H^{N}\sslash H^{0}\rightarrow H^{N}\sslash H$ is integral. It follows from [3, Theorem 6.3.3] applied to the quotient stack $[H^{N}/H^{0}]$ that $H^{N}\sslash H^{0}$ is of finite type over $S$. (If $S=\operatorname{Spec}R$, where $R$ is of finite type over a universally Japanese ring then the assertion also follows from [52, Theorem 2 (i)]. This setting is sufficient for our paper as we only need it over $\mathcal{O}$.) Hence, $H^{N}\sslash H^{0}\rightarrow H^{N}\sslash H$ is of finite type, and [54, Tag 01WJ] implies that $H^{N}\sslash H^{0}\rightarrow H^{N}\sslash H$ is finite.
@@
[…]
By 6.3 the right map is finite. Since the map $X^{\mathrm{gen},\tau}_{G}\to X^{\mathrm{ps}}_{\operatorname{GL}_{d}}\times G^{N}$ defined in 5.7 is a closed immersion, it follows from [52, Theorem 2 (ii)], that the induced map of GIT quotients, which is the top map in the diagram, is finite. It follows, that the left map in the diagram is finite. The same argument over $k$ proves the last assertion. ∎
@@ 6.1. Points in $X^{\mathrm{gen},\tau}_{G}$
    the orbit $G^{0}\cdot x$ is closed;
@@
    the orbit $G\cdot x$ is closed;
@@
It follows from (30) that (1) implies (2). If $G\cdot x$ is closed, then as it is $G^{0}$-invariant it will contain a closed $G^{0}$-orbit. We deduce from (30) that $G^{0}\cdot x$ is closed, so that (2) implies (1). We have established in Lemma 5.13 that $H$ is equal to the Zariski closure in $G(\kappa(x))$ of the subgroup generated by the $N$-tuple $(\rho_{x}(\gamma_{1}),\ldots,\rho_{x}(\gamma_{N}))$. The equivalence of parts (2), (3) has been established in [46, Theorem 16.4]. The equivalence of (3) and (4) follows from [5, Theorem 3.1], see the begining of [5, Section 6.3] for the non-connected case. ∎
@@
This follows from [52, Theorem 3]. ∎
@@
Let $x$ be a geometric point of $X^{\mathrm{gen},\tau}_{G}$. Then there is $z\in\overline{G^{0}\cdot x}$ such that the $G^{0}$-orbit of $\rho_{z}$ is equal to the $G$-semisimplification of $\rho_{x}$. Moreover, $G^{0}\cdot z$ is the unique closed $G^{0}$-orbit contained in $\overline{G^{0}\cdot x}$.
@@
Let $P_{\lambda}$ be a minimal R-parabolic containing $\rho_{x}(\Gamma_{F})$. Then
$\lim_{t\rightarrow 0}\lambda(t)\cdot(\rho_{x}(\gamma_{1}),\ldots,\rho_{x}(\gamma_{N}))$
exists and defines a point $z\in G^{N}(\kappa(x))$. Thus $z$ is contained in the closure of $G^{0}\cdot x$ in $G^{N}(\kappa(x))$. Since $X^{\mathrm{gen},\tau}_{G}$ is closed in $X^{\mathrm{ps}}_{\operatorname{GL}_{d}}\times G^{N}$ by 5.8, we deduce that $z\in X^{\mathrm{gen},\tau}_{G}(\kappa(x))$. Lemma 5.13 implies that the $G^{0}$-orbit of $\rho_{z}$ is the $G$-semisimplification of $\rho_{x}$, thus $\rho_{z}$ is $G$-completely reducible by Proposition 2.23. Thus the orbit $G\cdot z$ is closed by Proposition 6.7. Let $y$ be the image of $x$ in $X^{\mathrm{git},\tau}_{G}$. The fibre $(X^{\mathrm{gen},\tau}_{G})_{y}$ is a closed $G^{0}$-invariant subscheme $X^{\mathrm{gen},\tau}_{G}$ and thus it will contain $\overline{G^{0}\cdot x}$. The last assertion follows from Lemma 6.8. ∎
@@
By [52, Theorem 3 (ii)] $x$ and $x^{\prime}$ map to the same point in $X^{\mathrm{git},\tau}_{G}$ if and only if the closures of their $G^{0}$-orbits intersect non-trivially. If $\overline{G^{0}\cdot x}\cap\overline{G^{0}\cdot x^{\prime}}\neq\emptyset$ then it contains a point $z$ such that $G^{0}\cdot z$ is closed. It follows from Proposition 6.9 that the $G$-semisimplification of $\rho_{x}$ and of $\rho_{x^{\prime}}$ equal to the $G^{0}$-orbit of $\rho_{z}$. The other implication follows from directly from Proposition 6.9. ∎
@@ 6.2. Definition of $X^{\mathrm{gen},\tau}_{G,\overline{\rho}^{\mathrm{ss}}}$
6.4 implies that the map $R^{\mathrm{ps}}_{\operatorname{GL}_{d}}\rightarrow(A^{\mathrm{gen},\tau}_{G})^{G^{0}}$ is finite. Since $R^{\mathrm{ps}}_{\operatorname{GL}_{d}}$ is a complete local ring with residue field $k$, the Chinese remainder theorem implies that
@@
Since a connected component is a union of irreducible components, [8, Lemma 2.1] implies that the connected components of $X^{\mathrm{gen},\tau}_{G}\times_{X^{\mathrm{git},\tau}_{G}}\operatorname{Spec}R_{j}$ are $G^{0}$-invariant. It follows from [52, Theorem 3 (iii)] that if there were more than one then their images in $\operatorname{Spec}R_{j}$ under (33) would disconnect it. This is not possible as $R_{j}$ is a local ring. ∎
The representation $\overline{\rho}:\Gamma_{F}\rightarrow G(k)$ that we have fixed in Section 5 gives us a point $x_{0}\in X^{\mathrm{gen},\tau}_{G}(k)$. We denote the unique connected component of $X^{\mathrm{gen},\tau}_{G}$ on which $x_{0}$ lies on by $X^{\mathrm{gen},\tau}_{G,\overline{\rho}^{\mathrm{ss}}}$, where $\overline{\rho}^{\mathrm{ss}}$ is the $G$-semisimplification of the representation $\Gamma_{F}\overset{\overline{\rho}}{\longrightarrow}G(k)\hookrightarrow G(\overline{k})$. The justification for this notation is given in the lemma below. Let
@@
Let $z\in X^{\mathrm{gen},\tau}_{G,\overline{\rho}^{\mathrm{ss}}}(\kappa(x))$ be a point above $x_{0}$. It follows from Lemma 6.12 that $x$ and $z$ lie on the same connected component of $X^{\mathrm{gen},\tau}_{G}$ if and only if their images in $X^{\mathrm{git},\tau}_{G}$ lie on the same connected component of $X^{\mathrm{git},\tau}_{G}$. Since $z\in X^{\mathrm{gen},\tau}_{G,\overline{\rho}^{\mathrm{ss}}}$ and the residue field of $R^{\mathrm{git},\tau}_{G,\overline{\rho}^{\mathrm{ss}}}$ is $k$, $X^{\mathrm{git},\tau}_{G,\overline{\rho}^{\mathrm{ss}}}(\kappa(x))$ is a point, and thus this equivalent to $z$ and $x$ mapping to the same point in $X^{\mathrm{git},\tau}_{G}$. Corollary 6.10 implies that this is equivalent to $\rho_{x}$ and $\rho_{z}$ having the same $G$-semisimplification. ∎

### v2

@@ 5.1. Definition of $X^{\mathrm{gen}}_{\operatorname{GL}_{d}}$ and $X^{\mathrm{gen},\tau}_{G}$
If $L^{\prime}$ is a finite extension of $L$ with ring of integers $\mathcal{O}^{\prime}$ then it follows from the definition that $(X^{\mathrm{gen},\tau}_{G})_{\mathcal{O}^{\prime}}\cong X^{\mathrm{gen},\tau}_{G_{\mathcal{O}^{\prime}}}$. The main technical result of this paper is the computation of the dimension of $X^{\mathrm{gen},\tau}_{G}$. Since $\mathcal{O}^{\prime}$ is finite and flat over $\mathcal{O}$ the dimensions of $(X^{\mathrm{gen},\tau}_{G})_{\mathcal{O}^{\prime}}$ and $X^{\mathrm{gen},\tau}_{G}$ are the same and using Proposition 2.8 we may assume that $G$ is split, $G/G^{0}$ is constant and $G(\mathcal{O})\rightarrow(G/G^{0})(\mathcal{O})$ is surjective.
@@
Part (1) follows from Lemmas 4.15 and $5.$
@@
If $\mathfrak{p}$ is the maximal ideal of $R^{\mathrm{ps}}_{\operatorname{GL}_{d}}$ then $\kappa(\mathfrak{p})=k$ and we equip the algebraic closure $\overline{\kappa(\mathfrak{p})}$ with the discrete topology. If $\mathfrak{p}$ is a prime of $R^{\mathrm{ps}}_{\operatorname{GL}_{d}}$ with $\dim R^{\mathrm{ps}}_{\operatorname{GL}_{d}}/\mathfrak{p}=1$ then $\kappa(\mathfrak{p})$ is a local field and the valuation topology on $\kappa(\mathfrak{p})$ makes it into a topological $R^{\mathrm{ps}}_{\operatorname{GL}_{d}}$-algebra by [6, Lemma 3.17]. We equip the algebraic closure $\overline{\kappa(\mathfrak{p})}$ with the topology induced by a valuation, which extends the valuation on $\kappa(\mathfrak{p})$.
@@
Let $A^{\mathrm{gen}}_{\operatorname{GL}_{d}}\twoheadrightarrow A^{\mathrm{gen},\tau}_{G}$ be the quotient representing $X^{\mathrm{gen},\tau}_{G}$. We will denote by $\overline{X}^{\mathrm{gen},\tau}_{G}$ the special fibre of $X^{\mathrm{gen},\tau}_{G}$ over $\mathcal{O}$.
@@
Let $y:R^{\mathrm{ps}}_{\operatorname{GL}_{d}}\rightarrow\Omega$ be a ring homomorphism into an algebraically closed field $\Omega$ and let $x\in X^{\mathrm{gen},\tau}_{G}(\Omega)$ be a point above $y$. Let $H$ be the smallest Zariski closed subgroup of $G_{\Omega}$ containing $\rho_{x}(\gamma_{1}),\ldots,\rho_{x}(\gamma_{N})$, where $\rho_{x}$ is the specialisation at $x$ of the universal representation $\rho:\Gamma\rightarrow\operatorname{GL}_{d}(A^{\mathrm{gen}}_{\operatorname{GL}_{d}})$. Then $H$ is equal to the Zariski closure of $\rho_{x}(\Gamma)$ in $G_{\Omega}$.
@@ 5.2. Completion at maximal ideals and deformation problems
Let $Y\subset X^{\mathrm{gen},\tau}_{G}$ be the preimage of the closed point of $X^{\mathrm{ps}}_{\operatorname{GL}_{d}}$ and let $x$ be either a closed point of $Y$ or a closed point of $X^{\mathrm{gen},\tau}_{G}\setminus Y$ and let $y$ be its image in $\operatorname{Spec}R^{\mathrm{ps}}_{\operatorname{GL}_{d}}$. It follows from [6, Lemmas 3.17, 3.18] that $\kappa(x)$ is a finite extension of $\kappa(y)$ and there are the following possibilities:
@@
If $\rho_{A}\in D^{\square}_{\rho,G}(A)$ then [6, Proposition 3.33] implies that $A$ is naturally an $R^{\mathrm{ps}}_{\operatorname{GL}_{d}}$-algebra and $\tau\circ\rho\in X^{\mathrm{gen}}_{\operatorname{GL}_{d}}(A)$. Lemma 5.5 implies that there is a topologically finitely generated quotient $Q$ of $\Gamma$ such that any deformation of $\rho$ factor through $Q$. Hence, we may assume without loss of generality that $\gamma_{1},\ldots,\gamma_{N}$ are topological generators of $\Gamma$.
@@
By passing to the limit and using Lemma 5.15 with $A=R^{\square}_{\tau\circ\rho_{x},\operatorname{GL}_{d}}$ we obtain
@@
If $G=\operatorname{GL}_{d}$ and $\tau=\operatorname{id}$ then the proposition is [6, Proposition 3.33], so that $R^{\square}_{\tau\circ\rho_{x},\operatorname{GL}_{d}}$ is naturally isomorphic to the completion of the local ring of $\Lambda\otimes_{\mathcal{O}}A^{\mathrm{gen}}_{\operatorname{GL}_{d}}$ at $\mathfrak{q}$. Since $A^{\mathrm{gen},\tau}_{G}=A^{\mathrm{gen}}_{\operatorname{GL}_{d}}\otimes_{\mathcal{O}(\operatorname{GL}_{d}^{N})}\mathcal{O}(G^{N})=A^{\mathrm{gen}}_{\operatorname{GL}_{d}}/IA^{\mathrm{gen}}_{\operatorname{GL}_{d}}$ by definition, and $R^{\square}_{\tau\circ\rho_{x},\operatorname{GL}_{d}}$ is noetherian by Proposition 3.8, the assertion follows from Lemma 5.15 applied to $A=(\Lambda\otimes_{\mathcal{O}}A^{\mathrm{gen}}_{\operatorname{GL}_{d}})_{\mathfrak{q}}$. ∎
@@
The assertion follows from Proposition 5.16 and [6, Lemmas 3.36, 3.37]. ∎
@@ 6. GIT quotients
The closed immersion $\overline{X}^{\mathrm{gen},\tau}_{G}\hookrightarrow X^{\mathrm{gen},\tau}_{G}$ induces a morphism $\overline{X}^{\mathrm{git},\tau}_{G}\rightarrow(X^{\mathrm{git},\tau}_{G})_{k}$, which is finite by [57, Theorem 2] and induces a bijection on geometric points by [57, Theorem 3]. We deduce that this map is a homeomorphism.
@@
[30, Exp. V, Thm. 4.1 (ii)] applied with $X_{0}=H^{N}\sslash H^{0}$ and $X_{1}=(H/H^{0})\times_{S}X_{0}$, implies that the morphism $H^{N}\sslash H^{0}\rightarrow H^{N}\sslash H$ is integral. It follows from [1, Theorem 6.3.3] applied to the quotient stack $H^{N}/H^{0}]$ that $H^{N}\sslash H^{0}$ is of finite type over $S$. (If $S=\operatorname{Spec}R$, where $R$ is of finite type over a universally Japanese integral domain then the assertion also follows from [57, Theorem 2 (i)]. This setting is sufficient for our paper as we only need it over $\mathcal{O}$.) Hence, $H^{N}\sslash H^{0}\rightarrow H^{N}\sslash H$ is of finite type, and [59, Tag 01WJ] implies that $H^{N}\sslash H^{0}\rightarrow H^{N}\sslash H$ is finite.
@@
$X^{\mathrm{git},\tau}_{G}$$(X^{\mathrm{ps}}_{\operatorname{GL}_{d}}\times G^{N})\sslash G^{0}$$X^{\mathrm{ps}}_{\operatorname{GL}_{d}}$$(X^{\mathrm{ps}}_{\operatorname{GL}_{d}}\times\operatorname{GL}_{d}^{N})\sslash\operatorname{GL}_{d}$}
By 6.3 the right map is finite. Since the map $X^{\mathrm{gen},\tau}_{G}\to X^{\mathrm{ps}}_{\operatorname{GL}_{d}}\times G^{N}$ defined in 5.7 is a closed immersion, it follows from [57, Theorem 2 (ii)], that the induced map of GIT quotients, which is the top map in the diagram, is finite. It follows, that the left map in the diagram is finite. The same argument over $k$ proves the last assertion. ∎
@@ 6.1. Points in $X^{\mathrm{gen},\tau}_{G}$
    the orbit $G^{0}\cdot x$ is closed in its fibre over $X^{\mathrm{git},\tau}_{G}$;
@@
    the orbit $G\cdot x$ is closed in its fibre over $X^{\mathrm{git},\tau}_{G}$;
@@
It follows from (30) that (1) implies (2). If $G\cdot x$ is closed, then as it is $G^{0}$-invariant it will contain a closed $G^{0}$-orbit. We deduce from (30) that $G^{0}\cdot x$ is closed, so that (2) implies (1). We have established in Lemma 5.13 that $H$ is equal to the Zariski closure in $G(\kappa(x))$ of the subgroup generated by the $N$-tuple $(\rho_{x}(\gamma_{1}),\ldots,\rho_{x}(\gamma_{N}))$. The equivalence of parts (2), (3) has been established in [51, Theorem 16.4]. The equivalence of (3) and (4) follows from [3, Theorem 3.1], see the begining of [3, Section 6.3] for the non-connected case. ∎
@@
This follows from [57, Theorem 3]. ∎
@@
Let $x$ be a geometric point of $X^{\mathrm{gen},\tau}_{G}$. Then there is $z\in\overline{G^{0}\cdot x}$ such that the $G^{0}$-orbit of $\rho_{z}$ is equal to the $G$-semisimplification of $\rho_{x}$. Moreover, $G^{0}\cdot z$ is the unique closed $G^{0}$-orbit contained in $\overline{G^{0}\cdot x}$. Here the closures are taken in the fibre of $x$ over $X^{\mathrm{git},\tau}_{G}$.
@@
Let $z$ be a point in the closure of the $G^{0}$-orbit of $x$ in $G^{N}(\kappa(x))$, such that the $G^{0}$-orbit of $z$ is closed. Since $X^{\mathrm{gen},\tau}_{G}$ is closed in $X^{\mathrm{ps}}_{\operatorname{GL}_{d}}\times G^{N}$ by 5.8, we deduce that $z\in X^{\mathrm{gen},\tau}_{G}(\kappa(x))$. Lemma 5.13 implies that the $G^{0}$-orbit of $\rho_{z}$ is the $G$-semisimplification of $\rho_{x}$. Thus $\rho_{z}$ is $G$-completely reducible by Proposition 2.23. Thus the orbit $G\cdot z$ is closed by Proposition 6.7. Let $y$ be the image of $x$ in $X^{\mathrm{git},\tau}_{G}$. The fibre $(X^{\mathrm{gen},\tau}_{G})_{y}$ is a closed $G^{0}$-invariant subscheme $X^{\mathrm{gen},\tau}_{G}$ and thus it will contain $\overline{G^{0}\cdot x}$. The last assertion follows from Lemma 6.8. ∎
@@
Let $x$ and $x^{\prime}$ map to the same point in $X^{\mathrm{git},\tau}_{G}$. By 6.8 there is a unique closed $G^{0}$-orbit in the fibre of $y$, let $z\in X^{\mathrm{git},\tau}_{G}$ be a point in this orbit. It follows from Proposition 6.9 that the $G$-semisimplification of $\rho_{x}$ and of $\rho_{x^{\prime}}$ are equal to the $G^{0}$-orbit of $\rho_{z}$. The other implication follows from directly from Proposition 6.9. ∎
@@ 6.2. Definition of $X^{\mathrm{gen},\tau}_{G,\overline{\rho}^{\mathrm{ss}}}$
6.4 implies that the map $R^{\mathrm{ps}}_{\operatorname{GL}_{d}}\rightarrow(A^{\mathrm{gen},\tau}_{G})^{G^{0}}$ is finite. Since $R^{\mathrm{ps}}_{\operatorname{GL}_{d}}$ is a complete local ring with residue field $k$, part (1) of [59, Tag 04GH] implies that
@@
Since a connected component is a union of irreducible components, [6, Lemma 2.1] implies that the connected components of $X^{\mathrm{gen},\tau}_{G}\times_{X^{\mathrm{git},\tau}_{G}}\operatorname{Spec}R_{j}$ are $G^{0}$-invariant. It follows from [57, Theorem 3 (iii)] that if there were more than one then their images in $\operatorname{Spec}R_{j}$ under (33) would disconnect it. This is not possible as $R_{j}$ is a local ring. ∎
The representation $\overline{\rho}:\Gamma_{F}\rightarrow G(k)$ that we have fixed in Section 5 gives us a point $x_{0}\in X^{\mathrm{gen},\tau}_{G}(k)$. We denote the unique connected component of $X^{\mathrm{gen},\tau}_{G}$ which contains $x_{0}$ by $X^{\mathrm{gen},\tau}_{G,\overline{\rho}^{\mathrm{ss}}}$, where $\overline{\rho}^{\mathrm{ss}}$ is the $G$-semisimplification of the representation $\Gamma_{F}\overset{\overline{\rho}}{\longrightarrow}G(k)\hookrightarrow G(\overline{k})$. The justification for this notation is given in the lemma below. Let
@@
Let $z\in X^{\mathrm{gen},\tau}_{G,\overline{\rho}^{\mathrm{ss}}}(\kappa(x))$ be a point above $x_{0}$. It follows from Lemma 6.12 that $x$ and $z$ lie on the same connected component of $X^{\mathrm{gen},\tau}_{G}$ if and only if their images in $X^{\mathrm{git},\tau}_{G}$ lie on the same connected component of $X^{\mathrm{git},\tau}_{G}$. Since $z\in X^{\mathrm{gen},\tau}_{G,\overline{\rho}^{\mathrm{ss}}}$ and the residue field of $R^{\mathrm{git},\tau}_{G,\overline{\rho}^{\mathrm{ss}}}$ is $k$, $X^{\mathrm{git},\tau}_{G,\overline{\rho}^{\mathrm{ss}}}(\kappa(x))$ is a point, and thus this is equivalent to $z$ and $x$ mapping to the same point in $X^{\mathrm{git},\tau}_{G}$. Corollary 6.10 implies that this is equivalent to $\rho_{x}$ and $\rho_{z}$ having the same $G$-semisimplification. ∎